# Enhancing Automated Interpretability with Output-Centric Feature Descriptions

### v1

@@ 1 Introduction
To summarize, our work makes the following contributions: (a) we propose a two-faceted evaluation framework for feature descriptions, examining them through complementary input and output lenses (b) we highlight key drawbacks of using MaxAct, the common method used today in automated interpretability pipelines, to obtain feature descriptions in LLMs, (c) we propose output-centric methods to mitigate these limitations, (d) our experiments demonstrate the effectiveness of each approach and that their combination yields more faithful feature descriptions, (e) our analysis provides insights into the benefits in combining input- and output-centric methods. We release our code and generated feature descriptions at https://github.com/yoavgur/Feature-Descriptions.
@@ 2 Problem Setup
We focus on the problem of automatically describing atomic units of computation in LLMs called features. As the exact nature of features is a hotly debated topic, we adopt the general framework of Geiger et al. (2024a) which we limit to real-valued features. Any hidden vector $\mathbf{v}\in\mathbb{R}^{d}$ in the LLM can be transformed with an invertible featurizer $\mathcal{F}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ that maps the vector into a space of $k$ features. A single feature $f\in\mathbb{R}^{k}$ is simply a one-hot encoding which can be vectorized using $\mathbf{v}_{f}=\mathcal{F}^{-1}(f)$. This framework supports a variety of features, including neurons (axis-aligned dimensions) in MLPs Geva et al. (2022b), sets of orthogonal directions Geiger et al. (2024b); Huang et al. (2024); Park et al. (2024b), sparse linear features from SAEs Elhage et al. (2022); Templeton et al. (2024); Huben et al. (2024), or even non-linear features, e.g. “onion” representations with a magnitude-based features Csordás et al. (2024).
@@ 3 Evaluation of Feature Descriptions
Following Huang et al. (2023); Caden Juang et al. (2024), we evaluate how well the description captures the inputs triggering the feature. Given a feature $f$, we feed its description $s_{f}$ generated by some method into an LLM, which is tasked to generate two sets of $k$ examples each: activating and neutral. These examples are expected and not expected to activate $f$ according to $s_{f}$, respectively (see §A for examples and details regarding prompts). We then pass the generated examples through $\mathcal{M}$ and obtain $f$’s activation for each example, calculated as the max activation over all token positions in that example. We take the max over all token positions since it’s reasonable to expect $f$ to be activated highly even for just a single token, and not at all for the rest. Let $\bar{m}_{\text{activating}}$ and $\bar{m}_{\text{neutral}}$ be the mean activations obtained for the activating and neutral examples, respectively. The description $s_{f}$ is considered faithful if the mean activation for the activating examples exceeds that of the neutral examples, namely:
@@
To assess how faithful $s_{f}$ is with respect to $f$’s influence on the model’s outputs, we evaluate $s_{f}$ against outputs generated by $\mathcal{M}$ when steering $f$ versus when steering another feature $f^{\prime}$. Concretely, we feed $\mathcal{M}$ open-ended prompts, such as ‘‘<BOS> I think that’’ Chalnev et al. (2024), and let the model generate $n$ tokens three times – one time while amplifying $f$ and two other times by amplifying two different randoms features $f^{\prime}$ and $f^{\prime\prime}$. Amplification of a feature is done by clamping its activation to a high value $m$ Templeton et al. (2024). Since finding an effective yet not destructive amplification level is challenging Bhalla et al. (2024); Templeton et al. (2024), we run each input with varying levels of amplification while fixing the KL-divergence between the outputs of the steered model and the non-steered model Paulo et al. (2024), as calculated on a single next token prediction, averaged over all open ended prompts. This way we generate three sets of texts $\mathcal{T}_{f}$, $\mathcal{T}_{f^{\prime}}$ and $\mathcal{T}_{f^{\prime\prime}}$. Next, we feed $s_{f}$ concatenated with $\mathcal{T}_{f}$, $\mathcal{T}_{f^{\prime}}$ and $\mathcal{T}_{f^{\prime\prime}}$ to a judge LLM, and task it with indicating which of the three sets matches $s_{f}$. The description $s_{f}$ is faithful if the LLM selects $\mathcal{T}_{f}$. Namely, we evaluate how accurately the description captures the effect the feature has on the model’s output. For details, example generations and the exact prompts used, see §A.
@@ 4 Interpretability Methods
We describe the series of methods for automatically describing features in LLMs. These include the input-centric method prevalent today, two output-centric methods that describe a feature $f$ using its corresponding vector $\mathbf{v}_{f}$, and their ensembles.
@@
While both VocabProj and TokenChange are output-geared methods, VocabProj is correlative and TokenChange causally intervenes in the model’s generation.
@@ 5.1 Experimental Setting
We analyze both features learned through SAEs and neurons in MLP layers, covering four LLMs of different sizes and families: Gemma-2 2B Team et al. (2024b), Llama-3.1 8B and Llama-3.1 8B Instruct Dubey et al. (2024), and GPT-2 small Radford et al. (2019). For Gemma-2, Llama-3.1 and GPT-2 small, we evaluate descriptions of SAE features trained on residual stream and MLP layers: Gemma Scope 16K and 65K Lieberum et al. (2024), Llama Scope 32K He et al. (2024b), and OpenAI SAE 32K and 128K Gao et al. (2024). The activation function used by Gemma Scope is JumpReLU Rajamanoharan et al. (2024), while both Llama Scope and OpenAI SAE use TopK-ReLU Makhzani and Frey (2014). We randomly sample $n=40$ features per layer from every SAE, resulting in a total of 4,160 features for Gemma-2, 2,560 for Llama-3.1 and 1,440 for GPT-2 small. For Llama-3.1 Instruct we inspect a sample of $n=80$ MLP features per layer, with 2,560 features in total.
@@
We use the methods described in §4 and generate descriptions for each feature, using GPT-4o mini Hurst et al. (2024) as our explainer model to ensure consistency with descriptions from Neuronpedia Lin and Bloom (2023) and Transluce Choi et al. (2024). For MaxAct, we utilize the publicly available feature descriptions from these repositories. To validate these descriptions are comparable to the ones generated by us, we randomly sampled features and found their descriptions matched those we generate for MaxAct (see §B.3 for full analysis). When generating ensembles from raw data (Ensemble Raw), we rely on feature activation data from these same sources, using the top five activating sentences to keep in line with existing methods. Notably, Transluce generated descriptions for Llama-3.1 8B Instruct through a more complex process than MaxAct Choi et al. (2024), creating multiple descriptions from activating examples and selecting the best using simulation scoring Bills et al. (2023). For clarity, we refer to this method as MaxAct++ and generate the MaxAct descriptions for Llama-3.1 8B Instruct ourselves using the feature activation data from Transluce.
@@
For the input-based evaluation, we instruct Gemini 1.5 Pro Team et al. (2024a) to generate five activating and five neutral sentences with respect to a given feature description. For the output-based evaluation, we prompt the model with three different open-ended prompts, letting it generate up to 25 tokens while clamping the feature’s activation value to $m$ for all token positions. For each prompt, we run the model four times with increasing clamping values, making the generations progressively more affected by the feature’s output. The values were derived by fixing two target KL-divergence values $0.25$ and $0.5$, which provide two positive and two negative clamping values for $m$. We select these values to balance between generating text with sufficient feature effect and long or noisy text that is hard to evaluate. This process results in 12 text generations for each of the sets $\mathcal{T}_{\mathbf{v}_{f}}$, $\mathcal{T}_{\mathbf{v}_{f}^{\prime}}$, and $\mathcal{T}_{\mathbf{v}_{f}^{\prime\prime}}$. We then use Gemini 1.5 Pro Team et al. (2024a) as an LLM judge, as described in §3 (exact prompts in §A).
@@ 5.2 Results
Table 1 shows the results averaged across layers, and Figure 3 provides a breakdown for layer groups in Gemma-2 2B and Llama-3.1 8B Instruct. Similar trends are shown for Llama-3.1 and GPT-2 in §C.
@@
Table 1 shows that across all models and feature types, MaxAct outperforms VocabProj and TokenChange on the input-based evaluation and vice versa on the output-based evaluation, often by large margins of up to 15%-30%. This also holds for MaxAct++ on Llama-3.1 8B Instruct, demonstrating that input- and output-centric methods capture different feature information. Second, ensembling input- and output-centric methods boosts performance on both evaluations, with the ensembles combining all three methods consistently outperforming the single-methods. For instance, for Gemma-2 the ensembles yielded an improvement of 6%-10% over the next best single-method on both metrics. One exception to this trend is MaxAct++, which performs better than all other methods on the input metric, with Ensemble Raw (All) in close second. This is probably due to MaxAct++ being optimized for describing what activates a given feature. Overall, this input-output integration not only better describes the causal roles of features but also improves performance on the existing widely-used input-based evaluation.
@@
A major practical drawback of MaxAct is the computational cost required for comprehensively mapping the activating inputs of a feature. Considering the performance of VocabProj, TokenChange, and EnsembleR (VP+TC), we observe that (a) they typically outperform MaxAct on the output-based evaluation, which is crucial for assessing the description’s faithfulness to the feature’s causal effect and its usefulness for steering, and (b) they often perform only slightly worse on the input-based evaluation, e.g. there’s only a 3.5 point gap between Ensemble Raw (VP+TC) and MaxAct on residual stream SAE features in Gemma-2. These results suggest that VocabProj and TokenChange, which require only $\leq$2 inference passes, can often be a more efficient and sometimes higher-performing alternative to the widely-used MaxAct method.
@@ 6.1 Qualitative Analysis
We manually analyze the descriptions by MaxAct and VocabProj for a random sample of 100 features from Gemma Scope 16K, 50 for the MLP layers and 50 for the residual stream. We exclude TokenChange here as we noticed that the descriptions it produces are often similar to those by VocabProj (see examples in §E). In the analysis, we consider descriptions that pass both our input- and output-based evaluations. We observed 4 main types of relations between the descriptions:
@@
  • Composition: The input- and output- based descriptions refer to different aspects of the feature, while their composition provides a more holistic description of the feature.
@@
  • Different: The input- and output- based descriptions refer to different aspects of the feature, which share no clear relation between them.
@@ 6.2 Reviving Dead Features
We sampled 1,850 SAE features from Gemma-2 2B equally distributed across layers and types (MLP / residual) and classified as “dead” based on Neuronpedia. For each feature, we create a set of candidate prompts for activating it by: (a) using the feature descriptions by VocabProj and TokenChange and letting Gemini generate 150 sentences that are likely to activate the feature, and (b) gathering the tokens identified by VocabProj and TokenChange and constructing 1,450 sequences of different lengths that randomly combine these tokens. Both the top and bottom tokens obtained using these methods could potentially activate the feature, as they might relate to concepts that the feature promotes or suppresses. We then feed all the generated prompts into the the model and consider a feature as “revived” if any prompt successfully activated it. For implementation details, see §D.
@@ Results
The generated prompts successfully activated 9.1% (85) of MLP SAE features and 62% (491) of residual ones. In 12% (70) of cases, a feature was activated using an LLM-generated prompt, while 73% (423) were activated with a prompt composed of two tokens: ‘<BOS>’ and a sampled token. Moreover, the revived dead features can often be easily interpreted using VocabProj and TokenChange, while considered faithful based on our output-based metric (see examples in §D). Overall, this demonstrates that output-centric methods can address potential oversights that may arise from focusing solely on activating inputs.
@@ 7 Related Work
More broadly, our work relates to growing efforts in understanding features encoded in neurons and SAE features. These include steering Farrell et al. (2024); Chalnev et al. (2024); O’Brien et al. (2024b); Templeton et al. (2024), circuit discovery Marks et al. (2024); Makelov et al. (2024); Balcells et al. (2024), feature disentanglement Huang et al. (2024); Cohen et al. (2024) and benchmarks like SAEBench. [^1] https://www.neuronpedia.org/sae-bench/info However, evaluation of feature descriptions remains relatively underexplored. Rajamanoharan et al. (2024) evaluated latent interpretability for different SAE architectures using an input-centric approach which does not reflect downstream effect in model control. More recently, Paulo et al. (2024) have found negative correlation between multiple input-centric scoring methods and an intervention-based metric. Finally, Bhalla et al. (2024) concurrently evaluated feature descriptions in terms of their downstream effects on the model. However, they focus on evaluating methods for effectively steering models, as opposed to evaluating methods for generating descriptions.
@@ 8 Conclusion
Although we observe clear trends in the results, the output-based evaluation is fairly noisy. We address this by sampling large numbers of features and using multiple prompts in the evaluation, but future work could focus on reducing this noise further and making the evaluation more efficient. Additionally, we find that the output-centric methods and ensembles are sensitive to the choice of prompt. Since generating feature descriptions using these methods is non-trivial and often involves long texts (especially for the ensembles), improving explainer model prompts to extract relevant information could potentially enhance performance.
Regarding the methods evaluated, while we focused on efficient approaches that can automatically scale to millions of features, exploring other methods, such as patching-based methods, could be valuable. Lastly, the output-centric methods we propose are tied to the model’s vocabulary, which means they can only describe features that can be expressed with tokens from the vocabulary. These methods may struggle in describing features that are not easily or naturally expressed with words, such as positional features.

### v2

@@ 1 Introduction
To summarize, our work makes the following contributions: (a) we propose a two-faceted evaluation framework for feature descriptions, examining them through complementary input and output lenses (b) we highlight key drawbacks of using MaxAct, the common method used today in automated interpretability pipelines, to obtain feature descriptions in LLMs, (c) we propose output-centric methods to mitigate these limitations, (d) our experiments demonstrate the effectiveness of each approach and that their combination yields more faithful feature descriptions, (e) our analysis provides insights into the benefits in combining input- and output-centric methods. By producing more faithful and complete feature descriptions, our approach can enhance downstream applications such as model editing, machine unlearning, and circuit analysis (e.g., Wu et al., 2023; Farrell et al., 2024a; Marks et al., 2025). We release our code and generated feature descriptions at https://github.com/yoavgur/Feature-Descriptions.
@@ 2 Problem Setup
We focus on the problem of automatically describing atomic units of computation in LLMs called features. As the exact nature of features is a hotly debated topic, we adopt the general framework of Geiger et al. (2024a) which we limit to real-valued features. Let $\mathcal{M}$ be our target LLM. Any hidden vector $\mathbf{v}\in\mathbb{R}^{d}$ in $\mathcal{M}$ can be transformed with an invertible featurizer $\mathcal{F}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ that maps the vector into a space of $k$ features. A single feature $f\in\mathbb{R}^{k}$ is simply a one-hot encoding which can be vectorized using $\mathbf{v}_{f}=\mathcal{F}^{-1}(f)$. This framework supports a variety of features, including neurons (axis-aligned dimensions) in MLPs Geva et al. (2022b), sets of orthogonal directions Geiger et al. (2024b); Huang et al. (2024); Park et al. (2024b), sparse linear features from SAEs Bricken et al. (2023); Templeton et al. (2024); Huben et al. (2024), or even non-linear features, e.g. “onion” representations with a magnitude-based features Csordás et al. (2024).
@@ 3 Evaluation of Feature Descriptions
Following Huang et al. (2023); Caden Juang et al. (2024), we evaluate how well the description captures the inputs triggering the feature. Given a feature $f$, we feed its description $s_{f}$ generated by some method into an LLM, which is tasked to generate two sets of $k$ examples each: activating and neutral. These examples are expected and not expected to activate $f$ according to $s_{f}$, respectively (see §A for examples and details regarding prompts). We then pass the generated examples through $\mathcal{M}$ and obtain $f$’s activation for each example, calculated as the max activation over all token positions in that example. We take the max over all token positions since it’s reasonable to expect $f$ to be activated highly even for just a single token, and not at all for the rest, following prior work that treats strong localized activation as meaningful Bills et al. (2023); Choi et al. (2024); Paulo et al. (2024); Voita et al. (2024). Let $\bar{m}_{\text{activating}}$ and $\bar{m}_{\text{neutral}}$ be the mean activations obtained for the activating and neutral examples, respectively. The description $s_{f}$ is considered faithful if the mean activation for the activating examples exceeds that of the neutral examples, namely:
@@
To assess how faithful $s_{f}$ is with respect to $f$’s influence on the model’s outputs, we evaluate $s_{f}$ against outputs generated by $\mathcal{M}$ when steering $f$ versus when steering another feature $f^{\prime}$. Concretely, we feed $\mathcal{M}$ open-ended prompts, such as ‘‘<BOS> I think’’ Chalnev et al. (2024), and let the model generate $n$ tokens three times – one time while amplifying $f$ and two other times by amplifying two different random features $f^{\prime}$ and $f^{\prime\prime}$. Amplification of a feature is done by clamping its activation to a high value $m$ Templeton et al. (2024). Since finding an effective yet not destructive amplification level is challenging Bhalla et al. (2024); Templeton et al. (2024), we run each input with varying levels of amplification while fixing the KL-divergence between the outputs of the steered model and the non-steered model Paulo et al. (2024), as calculated on a single next token prediction, averaged over all open ended prompts. This way we generate three sets of texts $\mathcal{T}_{f}$, $\mathcal{T}_{f^{\prime}}$ and $\mathcal{T}_{f^{\prime\prime}}$. Next, we feed $s_{f}$ concatenated with $\mathcal{T}_{f}$, $\mathcal{T}_{f^{\prime}}$ and $\mathcal{T}_{f^{\prime\prime}}$ to a judge LLM (see justification in §E), and task it to indicate which of the three sets matches $s_{f}$. The description $s_{f}$ is faithful if the LLM selects $\mathcal{T}_{f}$. Namely, we evaluate how well the description captures the feature’s impact on the model’s output. For details, example generations and prompts used, see §A.
@@ 4 Interpretability Methods
We describe the methods used for automatically describing features in LLMs. These include the input-centric method prevalent today, two output-centric methods that describe a feature $f$ using its corresponding vector $\mathbf{v}_{f}$, and their ensembles.
@@
While both VocabProj and TokenChange are output-centric methods, VocabProj is correlative and TokenChange causally intervenes in the model’s generation.
@@ 5 Experiments
In this section, we evaluate the above methods on our input- and output-based evaluations. Additional human evaluations are reported in §E.
@@ 5.1 Experimental Setting
We analyze both features learned through SAEs and neurons in MLP layers, covering four LLMs of different sizes and families: Gemma-2 2B Team et al. (2024b), Llama-3.1 8B and Llama-3.1 8B Instruct Dubey et al. (2024), and GPT-2 small Radford et al. (2019). For Gemma-2, Llama-3.1 and GPT-2 small, we evaluate descriptions of SAE features trained on residual stream and MLP layers: Gemma Scope 16K and 65K Lieberum et al. (2024), Llama Scope 32K He et al. (2024b), and OpenAI SAE 32K and 128K Gao et al. (2024). The activation function used by Gemma Scope is JumpReLU Rajamanoharan et al. (2024), while both Llama Scope and OpenAI SAE use TopK-ReLU Makhzani and Frey (2014). We randomly sample $n=40$ features per layer from every SAE, resulting in a total of 4,160 features for Gemma-2, 2,560 for Llama-3.1 and 2,880 for GPT-2 small. For Llama-3.1 Instruct we inspect a sample of $n=80$ MLP features per layer, with 2,560 features in total.
@@
We use the methods described in §4 and generate descriptions for each feature, using GPT-4o mini Hurst et al. (2024) as our explainer model to ensure consistency with descriptions from Neuronpedia Lin and Bloom (2023) and Transluce Choi et al. (2024). For MaxAct, we utilize the publicly available feature descriptions from these repositories. To validate these descriptions are comparable to those generated by us, we sampled 1,080 features and found their descriptions match those we generate for MaxAct (see §B.3).
When generating ensembles from raw data (Ensemble Raw), we rely on feature activation data from these same sources, using the top five activating sentences to keep in line with existing methods. Notably, Transluce generated descriptions for Llama-3.1 8B Instruct through a more complex process than MaxAct Choi et al. (2024), creating multiple descriptions from activating examples and selecting the best one using simulation scoring Bills et al. (2023). For clarity, we refer to this method as MaxAct++ and generate the MaxAct descriptions for Llama-3.1 8B Instruct ourselves using the feature activation data from Transluce.
@@
For the input-based evaluation, we instruct Gemini 1.5 Pro Team et al. (2024a) to generate five activating and five neutral sentences with respect to a given feature description. For the output-based evaluation, we prompt the model with three open-ended prompts, letting it generate up to 25 tokens while clamping the feature’s activation value to $m$ for all token positions. For each prompt, we run the model four times with increasing clamping values, making the generations progressively more affected by the feature’s output. This process results in 12 text generations for each of the sets $\mathcal{T}_{\mathbf{v}_{f}}$, $\mathcal{T}_{\mathbf{v}_{f}^{\prime}}$, and $\mathcal{T}_{\mathbf{v}_{f}^{\prime\prime}}$, which we provide to GPT-4o mini Hurst et al. (2024) as a judge (see §A for more details and exact prompts). We select this model to minimize costs, given the lengthy prompts induced by the text sets.
@@ 5.2 Results
Table  1 shows the results averaged across layers, and Figure 3 provides a breakdown for layer groups for features from Gemma-2 and both Llama-3.1 models. Similar trends are shown for all other features in §C.
@@
Table 1 shows that across all models and feature types, MaxAct outperforms VocabProj and TokenChange on the input-based evaluation and vice versa on the output-based evaluation, often by large margins of up to 15%-30%. This also holds for MaxAct++ on Llama-3.1 8B Instruct, demonstrating that input- and output-centric methods capture different feature information. Second, ensembling input- and output-centric methods boosts performance on both evaluations, with the ensembles combining all three methods consistently outperforming the single-methods. For instance, for Gemma-2 the ensembles yielded an improvement of 6%-10% over the next best single-method on both metrics. One exception to this trend is MaxAct++, which performs better than all other methods on the input metric, with Ensemble Raw in close second. This is probably due to MaxAct++ being optimized for describing what activates a given feature. Overall, this input-output integration not only better describes the causal roles of features but also improves performance on the widely-used input-based evaluation.
@@
A major practical drawback of MaxAct is the computational cost required for comprehensively mapping the activating inputs of a feature. Considering the performance of VocabProj, TokenChange, and EnsembleR (VP+TC), we observe that (a) they typically outperform MaxAct on the output-based evaluation, which is crucial for assessing the description’s faithfulness to the feature’s causal effect and its usefulness for steering, and (b) they often perform only slightly worse on the input-based evaluation, e.g. there’s only a 3.5 point gap between Ensemble Raw (VP+TC) and MaxAct on residual stream SAE features in Gemma-2. These results suggest that VocabProj and TokenChange, which require only $\leq$2 inference passes, can often be a more efficient and sometimes higher-performing alternative to the widely-used MaxAct method. An analysis of the computational costs is in §D.
@@ 6.1 Qualitative Analysis
We manually analyze the descriptions by MaxAct and VocabProj for a random sample of 100 features from Gemma Scope 16K, 50 for the MLP layers and 50 for the residual stream. We exclude TokenChange here as we noticed that the descriptions it produces are often similar to those by VocabProj (see examples in §G). In the analysis, we consider descriptions that pass both our input- and output-based evaluations. We observed 4 main types of relations between the descriptions:
@@
  • Composition: The input- and output-centric descriptions refer to different aspects of the feature, while their composition provides a more holistic description of the feature.
@@
  • Different: The input- and output-centric descriptions refer to different aspects of the feature, which share no clear relation between them.
@@ 6.2 Reviving Dead Features
We sampled 1,850 SAE features from Gemma-2 2B equally distributed across layers and types (MLP / residual) and classified as “dead” based on Neuronpedia. For each feature, we create a set of candidate prompts for activating it by: (a) using the feature descriptions by VocabProj and TokenChange and letting Gemini generate 150 sentences that are likely to activate the feature, and (b) gathering the tokens identified by VocabProj and TokenChange and constructing 1,450 sequences of different lengths that randomly combine these tokens. Both the top and bottom tokens obtained using these methods could potentially activate the feature, as they might relate to concepts that the feature promotes or suppresses. We then feed all the generated prompts into the model and consider a feature as “revived” if any prompt successfully activated it. For implementation details, see §F.
@@ Results
The generated prompts successfully activated 9.1% (85) of MLP SAE features and 62% (491) of residual ones. In 12% (70) of cases, a feature was activated using an LLM-generated prompt, while 73% (423) were activated with a prompt composed of two tokens: ‘<BOS>’ and a sampled token. Moreover, the revived dead features can often be easily interpreted using VocabProj and TokenChange, while considered faithful based on our output-based metric (see examples in §F). Overall, this demonstrates that output-centric methods can address potential oversights that may arise from focusing solely on activating inputs.
@@ 7 Related Work
More broadly, our work relates to growing efforts in understanding features encoded in neurons and SAE features. These include steering Farrell et al. (2024b); Chalnev et al. (2024); O’Brien et al. (2024b); Templeton et al. (2024), circuit discovery Marks et al. (2024); Makelov et al. (2024); Balcells et al. (2024), feature disentanglement Huang et al. (2024); Cohen et al. (2024) and benchmarks like SAEBench. [^1] https://www.neuronpedia.org/sae-bench/info However, evaluation of feature descriptions remains relatively underexplored. Rajamanoharan et al. (2024) evaluated latent interpretability for different SAE architectures using an input-centric approach which does not reflect downstream effect in model control. More recently, Paulo et al. (2024) have found negative correlation between multiple input-centric scoring methods and an intervention-based metric. Finally, Bhalla et al. (2024) concurrently evaluated feature descriptions in terms of their downstream effects on the model. However, they focus on evaluating methods for effectively steering models, as opposed to evaluating methods for generating descriptions.
@@ 8 Conclusion
Although we observe clear trends in the results, the output-based evaluation is fairly noisy. We address this by sampling large numbers of features and using multiple prompts in the evaluation, but future work could focus on reducing this noise further and making the evaluation more efficient. Additionally, we find that the output-centric methods and ensembles are sensitive to the choice of prompt. Since generating feature descriptions using these methods is non-trivial and often involves long texts (especially for the ensembles), improving explainer model prompts to extract relevant information could potentially enhance performance. We also note that our input-based evaluation uses a binary threshold, which may oversimplify feature behavior. Nonetheless, it enabled us to efficiently identify trends across models and methods, and we leave refining this evaluation to future work.
Regarding the methods evaluated, while we focused on efficient approaches that can automatically scale to millions of features, exploring other methods, such as patching-based methods, could be valuable. Lastly, the output-centric methods we propose are tied to the model’s vocabulary, which means they can only describe features that can be expressed with tokens from the vocabulary. These methods may struggle in describing features that are not easily or naturally expressed with words, such as positional features. For simplicity, we did not differentiate between whether concepts were being suppressed or promoted by a feature.